# Detection of MEMS Acoustics via Scanning Tunneling Microscopy

### v1

@@ I Introduction
In recent decades, nano-membranes have facilitated breakthroughs in quantum science and precision sensing. Their thin geometries create minimal physical connection to external environments, which, combined with high tensile stresses, allows for dissipation dilution that achieves state-of-the-art mechanical isolation from surrounding environments. This isolation enables the probing of macroscopic quantum phenomena present in the membranes. Nano-membranes are widely used in quantum-limited sensing, where their relatively large surface areas make them ideal for sensing macroscopic forces such as pressure, acceleration, and Casimir effects [Xu2025arXiv].
Common interrogation methods for membrane acoustics includes optics [Thompson2008Nature, norte2016, Reinhardt2016PRX, tsaturyan2017ultracoherent] and electronics [andrews2014bidirectional, malcovati2018evolution, zawawi2020review]. However, these methods require strong interaction with the membrane’s, potentially introducing noise or perturbations via heating and charging, obscuring properties such as superconductivity or ferromagnetism. Recently, techniques using scanning force microscopy on graphene membranes has been presented [Halg2020PRA]. While less invasive, such techniques still utilize lasers for the excitation and measurement of membrane resonances. Scanning microwave microscopy has been used to image acoustic modes of small silicon nitride membranes with microwave tips which are 50-550nm away from the membrane [xu2024imaging]. In contrast, here we use a scanning tunneling microscope (STM) tip to both excite and measure the resonances of (super)conducting membranes. The STM tip is uniquely capable of measuring the local density of states on surfaces with sub-nanometer precision via small tunneling currents. We show the capability to excite the membrane while simultaneously probing its oscillation amplitude using STM, achieving subatomic spatial precision with minimal invasiveness.
It is important to note that since 2012 there have been a number of studies [uder2018low, breitwieser2017, alyobi2020voltage, zan2012scanning, klimov2012electromechanical, xu2012atomic, zhao2013fabrication, breitwieser2014flipping, schoelz2015graphene, uder2017convenient] using scanning tunneling microscopy to approach free-standing graphene membranes of dimensions ranging from a few microns [uder2018low, alyobi2020voltage] to tens of microns [breitwieser2017]. These studies focused on manipulating wrinkling on these highly compliant membranes and manipulating their shapes, rather than engaging with their resonant acoustic modes. In contrast, our work enables the driving and measurement of acoustic modes in membranes that are approximately 100,000 times more massive than any other object manipulated with scanning tunneling microscopy to date. In doing so, we connect mesoscopic mechanical motion to atomic-scale quantum tunneling, establishing a platform where macroscopic forces can be sensed through what is in essence a quantum process.
@@ II Experimental Set-up
Our membrane samples consist of silicon substrates with deposited metal layers, which are patterned with holes to remove the underlying substrate, leaving behind suspended, high–aspect-ratio membranes. We focus on two types: a more compliant dielectric SiN membrane coated with superconducting NbTiN  [thoen2016superconducting] and a stiffer gold-coated membrane. We attribute the difference in compliance to the gold being deposited via e-beam evaporation, whereas the NbTiN is grown by atomic layer deposition (see Methods section for more details). We find a Q-factor of over $10^{6}$ for the NbTiN coated membrane, see Supplementary Note 2.
@@
where the AC component with amplitude $V_{\mathrm{drive}}$ actuates the membrane motion while the DC offset $V_{\mathrm{DC}}$ stabilizes the tunneling current used by the STM feedback loop. The membrane oscillations are not observed in real time because the $1.1$ kHz bandwidth of the current pre-amplifier is far below the membrane’s mechanical resonance frequencies, which in this study lie between $100$ kHz and $1$ MHz. Instead, we detect the motion indirectly through three different modalities, explained below. Using a nested feedback loop to maintain a stable tip current, we also maintain lateral precision down to the picometer scale. As shown later (Fig. 4a,b), this positional stability is essential for reproducibility, since even subnanometer-scale lateral shifts in tip position alter the local tip–membrane coupling strength. This requirement arises from the STM’s exceptionally small effective interrogation area, set by the short-range nature of electron tunneling and near-field tip–sample forces, which is expected to be on the order of $S_{\rm S}\approx 10^{-16}$ m2 (see Sec. 11.3 of Ref. [Israelachvili2011]). As illustrated in Fig. 1b, this nanoscale interrogation window distinguishes our approach from other membrane readout techniques such as optical interferometry (typical spot $S_{\rm L}\approx 10^{-10}$ m2 [Thompson2008Nature, Reinhardt2016PRX, norte2016]), capacitive sensing ($S_{\rm C}\approx 10^{-8}$ m2 [Faust2012PRL, andrews2014bidirectional, zawawi2020review]), or piezoelectric detection ($S_{\rm P}\approx 10^{-6}$ m2 [Mahboob2008APL, Elfrink2009JMM]). The STM thus accesses a spatially localized regime of membrane dynamics unattainable by macroscopic probe geometries. Furthermore, STM can provide unique insight into the electronic properties of the membrane (see Suppl. Note 3).
@@
In the case of Homodyne Detection, the response corresponds to the difference in tunneling current between measurements with and without the oscillating voltage. This is achieved by chopping the excitation at a reference frequency $f_{\rm LI}$ well below the pre-amplifier cutoff and then demodulating the resulting signal at this same frequency. This yields a lock-in signal that carries phase information (see Supplementary Note 4 for more details), and is inspired by recent ESR-STM experiments, operating at GHz frequencies [Hwang2022RevSciIns]. This lock-in signal scales approximately linearly with the driving amplitude and therefore imposes a lower limit to the usable drive power. Homodyne Detection, by virtue of multiplying an oscillating voltage with an oscillating conductance (through the oscillating tip–membrane distance) at equal frequency, produces a phase-sensitive DC signal scaling as $\cos({\vartheta_{\rm elec}-\vartheta_{\rm mech}})$, where $\vartheta_{\rm mech}$is the phase of the membrane’s mechanical motion and $\vartheta_{\rm elec}$ is the phase of the oscillating electrical signal. This DC component is further enhanced by the exponential dependence of the tunneling conductance on the tip–membrane distance (see Supplementary Note 4). When combined with lock-in detection to reject out-of-band noise, this signal allows even sub-picometer membrane oscillations to be measured. Figure 1c shows the lock-in signal, after rejecting the capacitive background via phase adjustment, measured on a stiff gold-coated membrane.
@@ 9 Sweeps $f_{\rm drive}$ / $f_{0}$ $f_{\rm drive}$ $f_{\rm drive}$ $f_{0}$
Table 1 summarizes the main differences in applicability between the three detection modalities. Homodyne Detection is preferred when phase information is required, for instance, to determine whether the oscillating electric field locally pushes or pulls the membrane, to verify that the electrical coupling between tip and membrane is stable, or to identify whether the response remains in the linear regime. Feedback Resonance is most suitable when measurements need to be performed rapidly, such as during limited cryogenic hold times, or when tracking the resonance frequency efficiently as it shifts under changing external conditions. For example, in our related work where we measure the Casimir force across the membrane’s superconducting transition [Xu2025arXiv], we required low driving powers and a fast and reliable measurement modality to overcome our limited Liquid Helium hold time at elevated temperatures, thus making Feedback Resonance a natural choice. Z-Sweep Resonance, on the other hand, is the method of choice when one aims to minimize the influence of the STM tip on the membrane altogether and measure its intrinsic mechanical response in the least perturbative way.
@@ IV Electrostatic Forces
The last term in Eq.  2 is determined by the electrostatic force and scales quadratically with $V_{\rm DC}$. Varying this potential allows extraction of the local contact potential differential (LCPD) $V_{\rm LCPD}$, using either the z-sweep and feedback resonance modalities. This approach is analogous to Kelvin-probe measurements in AFM [Bettac2009Nanotech], where $V_{\rm DC}$ compensates the chemical potential offset between tip and membrane. Figure 3a shows z-sweep resonance curves at different $V_{\rm DC}$. As $|V_{\rm DC}|$ increases, the onset of current rise $z_{\rm rise}$ shifts to larger tip-membrane distances, indating that the electrostatic attraction modifies the resonance frequency $f_{0}$. Converting these $z_{\rm rise}$ using a height-to-frequency mapping (inset) yields a parabolic dependence of frequency detuning on bias (panel c), characteristic of LCPD behavior. The mapping in the inset was made by finding $z_{\rm rise}$ in repeated Z-Sweep Resonance measurements at different $f_{\rm drive}$.
@@ VII Conclusions
Future work could enhance frequency resolution further, while the method’s sensitivity to large-scale forces, such as acceleration, pressure, and surface forces, positions it as a promising tool in quantum sensing applications. Examples include detecting Casimir forces [Xu2025arXiv] or monitoring nanoscale growth in real time. Furthermore, by combining this technique with higher Q-factor membranes, it may be possible to explore more complex macroscopic quantum phenomena, making STM-based nanomechanical sensing a versatile platform for both fundamental research and applied science.
@@ VIII.2 Mathematical Model
After extracting the coefficients from a fit, we can then simulate the following equation of motion using numerical continuation [doedel2007auto] in order to obtain the nonlinear steady-state response of the driven system:
@@ VIII.3 Membrane Navigation System
We did not optimize the tip apex for properly resolving the small markers. We expect STM tips that are prepared more carefully, such as through a combination of Field Ion Microscopy (FIM) and Field Emission Microscopy (FEM) to benefit more from this navigation system [Gomer1961Harvard, Muller1969Elsevier].
@@ VIII.4 Sample Preparation
Detection of MEMS Acoustics via Scanning Tunneling Microscopy (Supplementary)
R. J. G. Elbertse§
M. Xu§

### v2

@@ Abstract
Corresponding authors: a.f.otte@tudelft.nl, r.a.norte@tudelft.nl
Corresponding author:
@@ I Introduction
In recent decades, nano-membranes have facilitated breakthroughs in quantum science and precision sensing. Their thin geometries create minimal physical connection to external environments, which, combined with high tensile stresses, allows for dissipation dilution that achieves state-of-the-art mechanical isolation from surrounding environments. This isolation enables the probing of macroscopic quantum phenomena present in the membranes. Nano-membranes are widely used in quantum-limited sensing, where their relatively large surface areas make them ideal for sensing macroscopic forces such as pressure, acceleration, and Casimir effects [26].
Common interrogation methods for membrane acoustics includes optics [21, 17, 18, 22] and electronics [2, 15, 29]. However, these methods require strong interaction with the membrane’s, potentially introducing noise or perturbations via heating and charging, obscuring properties such as superconductivity or ferromagnetism. Recently, techniques using scanning force microscopy on graphene membranes has been presented [10]. While less invasive, such techniques still utilize lasers for the excitation and measurement of membrane resonances. Scanning microwave microscopy has been used to image acoustic modes of small silicon nitride membranes with microwave tips which are 50-550nm away from the membrane [25]. In contrast, here we use a scanning tunneling microscope (STM) tip to both excite and measure the resonances of (super)conducting membranes. The STM tip is uniquely capable of measuring the local density of states on surfaces with sub-nanometer precision via small tunneling currents. We show the capability to excite the membrane while simultaneously probing its oscillation amplitude using STM, achieving subatomic spatial precision with minimal invasiveness.
It is important to note that since 2012 there have been a number of studies [24, 4, 1, 28, 13, 27, 30, 5, 19, 23] using scanning tunneling microscopy to approach free-standing graphene membranes of dimensions ranging from a few microns [24, 1] to tens of microns [4]. These studies focused on manipulating wrinkling on these highly compliant membranes and manipulating their shapes, rather than engaging with their resonant acoustic modes. In contrast, our work enables the driving and measurement of acoustic modes in membranes that are approximately 100,000 times more massive than any other object manipulated with scanning tunneling microscopy to date. In doing so, we connect mesoscopic mechanical motion to atomic-scale quantum tunneling, establishing a platform where macroscopic forces can be sensed through what is in essence a quantum process.
@@ II Experimental Set-up
Our membrane samples consist of silicon substrates with deposited metal layers, which are patterned with holes to remove the underlying substrate, leaving behind suspended, high–aspect-ratio membranes. We focus on two types: a more compliant dielectric SiN membrane coated with superconducting NbTiN  [20] and a stiffer gold-coated membrane. We attribute the difference in compliance to the gold being deposited via e-beam evaporation, whereas the NbTiN is grown by atomic layer deposition (see Methods section for more details). We find a Q-factor of over $10^{6}$ for the NbTiN coated membrane, see Supplementary Note 2.
@@
where the AC component with amplitude $V_{\mathrm{drive}}$ actuates the membrane motion while the DC offset $V_{\mathrm{DC}}$ stabilizes the tunneling current used by the STM feedback loop. The membrane oscillations are not observed in real time because the $1.1$ kHz bandwidth of the current pre-amplifier is far below the membrane’s mechanical resonance frequencies, which in this study lie between $100$ kHz and $1$ MHz. Instead, we detect the motion indirectly through three different modalities, explained below. Using a nested feedback loop to maintain a stable tip current, we also maintain lateral precision down to the picometer scale. As shown later (Fig. 4a,b), this positional stability is essential for reproducibility, since even subnanometer-scale lateral shifts in tip position alter the local tip–membrane coupling strength. This requirement arises from the STM’s exceptionally small effective interrogation area, set by the short-range nature of electron tunneling and near-field tip–sample forces, which is expected to be on the order of $S_{\rm S}\approx 10^{-16}$ m2 (see Sec. 11.3 of Ref. [12]). As illustrated in Fig. 1b, this nanoscale interrogation window distinguishes our approach from other membrane readout techniques such as optical interferometry (typical spot $S_{\rm L}\approx 10^{-10}$ m2 [21, 18, 17]), capacitive sensing ($S_{\rm C}\approx 10^{-8}$ m2 [8, 2, 29]), or piezoelectric detection ($S_{\rm P}\approx 10^{-6}$ m2 [14, 7]). The STM thus accesses a spatially localized regime of membrane dynamics unattainable by macroscopic probe geometries. Furthermore, STM can provide unique insight into the electronic properties of the membrane (see Suppl. Note 3).
@@
In the case of Homodyne Detection, the response corresponds to the difference in tunneling current between measurements with and without the oscillating voltage. This is achieved by chopping the excitation at a reference frequency $f_{\rm LI}$ well below the pre-amplifier cutoff and then demodulating the resulting signal at this same frequency. This yields a lock-in signal that carries phase information (see Supplementary Note 4 for more details), and is inspired by recent ESR-STM experiments, operating at GHz frequencies [11]. This lock-in signal scales approximately linearly with the driving amplitude and therefore imposes a lower limit to the usable drive power. Homodyne Detection, by virtue of multiplying an oscillating voltage with an oscillating conductance (through the oscillating tip–membrane distance) at equal frequency, produces a phase-sensitive DC signal scaling as $\cos({\vartheta_{\rm elec}-\vartheta_{\rm mech}})$, where $\vartheta_{\rm mech}$is the phase of the membrane’s mechanical motion and $\vartheta_{\rm elec}$ is the phase of the oscillating electrical signal. This DC component is further enhanced by the exponential dependence of the tunneling conductance on the tip–membrane distance (see Supplementary Note 4). When combined with lock-in detection to reject out-of-band noise, this signal allows even sub-picometer membrane oscillations to be measured. Figure 1c shows the lock-in signal, after rejecting the capacitive background via phase adjustment, measured on a stiff gold-coated membrane.
@@ 9 Sweeps $f_{\rm drive}$ / $f_{0}$ $f_{\rm drive}$ $f_{\rm drive}$ $f_{0}$
Table 1 summarizes the main differences in applicability between the three detection modalities. Homodyne Detection is preferred when phase information is required, for instance, to determine whether the oscillating electric field locally pushes or pulls the membrane, to verify that the electrical coupling between tip and membrane is stable, or to identify whether the response remains in the linear regime. Feedback Resonance is most suitable when measurements need to be performed rapidly, such as during limited cryogenic hold times, or when tracking the resonance frequency efficiently as it shifts under changing external conditions. For example, in our related work where we measure the Casimir force across the membrane’s superconducting transition [26], we required low driving powers and a fast and reliable measurement modality to overcome our limited Liquid Helium hold time at elevated temperatures, thus making Feedback Resonance a natural choice. Z-Sweep Resonance, on the other hand, is the method of choice when one aims to minimize the influence of the STM tip on the membrane altogether and measure its intrinsic mechanical response in the least perturbative way.
@@ IV Electrostatic Forces
The last term in Eq.  2 is determined by the electrostatic force and scales quadratically with $V_{\rm DC}$. Varying this potential allows extraction of the local contact potential differential (LCPD) $V_{\rm LCPD}$, using either the z-sweep and feedback resonance modalities. This approach is analogous to Kelvin-probe measurements in AFM [3], where $V_{\rm DC}$ compensates the chemical potential offset between tip and membrane. Figure 3a shows z-sweep resonance curves at different $V_{\rm DC}$. As $|V_{\rm DC}|$ increases, the onset of current rise $z_{\rm rise}$ shifts to larger tip-membrane distances, indating that the electrostatic attraction modifies the resonance frequency $f_{0}$. Converting these $z_{\rm rise}$ using a height-to-frequency mapping (inset) yields a parabolic dependence of frequency detuning on bias (panel c), characteristic of LCPD behavior. The mapping in the inset was made by finding $z_{\rm rise}$ in repeated Z-Sweep Resonance measurements at different $f_{\rm drive}$.
@@ VII Conclusions
Future work could enhance frequency resolution further, while the method’s sensitivity to large-scale forces, such as acceleration, pressure, and surface forces, positions it as a promising tool in quantum sensing applications. Examples include detecting Casimir forces [26] or monitoring nanoscale growth in real time. Furthermore, by combining this technique with higher Q-factor membranes, it may be possible to explore more complex macroscopic quantum phenomena, making STM-based nanomechanical sensing a versatile platform for both fundamental research and applied science.
@@ VIII.2 Mathematical Model
After extracting the coefficients from a fit, we can then simulate the following equation of motion using numerical continuation [6] in order to obtain the nonlinear steady-state response of the driven system:
@@ VIII.3 Membrane Navigation System
We did not optimize the tip apex for properly resolving the small markers. We expect STM tips that are prepared more carefully, such as through a combination of Field Ion Microscopy (FIM) and Field Emission Microscopy (FEM) to benefit more from this navigation system [9, 16].